OpenStudy (anonymous):

Evaluate C(6, 6). Using combination formula.

ganeshie8 (ganeshie8):

it is asking to use combination formula

ganeshie8 (ganeshie8):

hmm you must be having the formula in your notes, can you check

OpenStudy (anonymous):

C(n,r)=\[\frac{ n! }{r! * (n-4)}\]

OpenStudy (anonymous):

but it's the same number, do I use the same formula?

ganeshie8 (ganeshie8):

C(n,r) = \(\frac{ n! }{r! * (n-r)} \)

ganeshie8 (ganeshie8):

yes, we use the formula cuz question is asking us to use.

ganeshie8 (ganeshie8):

C(6, 6) \(\large \frac{6!}{6!(6-6)!} \) \(\large \frac{\cancel{6!}}{\cancel{6!}(0)!} \)

OpenStudy (anonymous):

so then \[\frac{ 6! }{ 6!*(6-6)! }\]

OpenStudy (anonymous):

oh, lol

ganeshie8 (ganeshie8):

:) whats the answer

OpenStudy (anonymous):

1

ganeshie8 (ganeshie8):

perfect ! we can confirm it by seeing that, u can pick 6 things from 6 things, in only 1 way

OpenStudy (anonymous):

thanks :D

ganeshie8 (ganeshie8):

yw :)